WHAT IS THE POSITIVE SLOPE OF A RAMP WITH A VERTICAL RISE OF 24 FEET AND A HORIZONTAL RUN OF 40 FEET ?
Answer by rfer(16322)

24/40=3/5 or 0.6=60%